OpenStudy (anonymous):

SETTING UP PROBLEM Two planes make a 1500 mile flight, one flying 50 miles per hour faster than the other. The quicker plane makes the trip 1 hour faster. How long did it take the slower plane to complete the flight

OpenStudy (anonymous):

someone help

OpenStudy (anonymous):

jus trying

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let slower plane speed =x then faster plane speed=x+50 distance=1500 time=distance/velocity so time taken for slower plane=1500/x time taken for slower plane=1500/x+50 difference between them is 1hr 1500/x+50 -1500/x==1 solve for x and fnd the time using 1500/x+50

OpenStudy (anonymous):

speed = distanve * time let speed of first plane be v, d = 1500, t = time taken speed of the second plane is 50 faster (v+50) and time taken is 1 slower = d-1 First plane: v = 1500 / t Second plane: v + 50 = 1500 / (t - 1) solving these two equations will give you the answer

OpenStudy (anonymous):

its 6 hrs

OpenStudy (anonymous):

THANKS EVERYONE!

OpenStudy (anonymous):

let the slower plane took t hours. so faster plane took (t-1)hours speed of slower plane=(1500/t) miles/hr speed of faster plane={1500/(t-1)} miles/hr acc. to ques., {1500/(t-1)}=(1500/t)+50 solve for t which is ur answer.

OpenStudy (anonymous):

t=1500/v v+50=1500/(t-1) solve t